Maria has invested a certain amount of money in stock and bond. The annual return on these investments is 570$. If the stock produces 150$ more in annual return then bond, how much money does Maria receive per year from each type of investment?
Answer by TimothyLamb(4379)

x = return from stocks
y = return from bonds
---
x + y = 570
x = y + 150
---
put the system of linear equations into standard form
---
x + y = 570
x - y = 150
---
copy and paste the above standard form linear equations in to this solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
solution:
x = return from stocks = 360
y = return from bonds = 210
---
Free algebra tutoring live chat:
https://sooeet.com/chat.php?gn=algebra
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations with quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php
---